Baker Ivanov organizes a pastry shop. He rents premises for 10 thousand UAH. per year, invites a pastry chef, whose salary is 24 thousand UAH. per year, and buys raw materials for 40 thousand UAH. in year. The production equipment owned by Ivanov is estimated at UAH 80,000. This amount could bring him an annual income of UAH 8 thousand. The annual depreciation of the equipment is 10 thousand UAH.

Prior to that, Ivanov, working at the bakery, received a salary of 30 thousand hryvnyas. in year. Ivanov knows that he has entrepreneurial abilities and estimates them at UAH 6,000.

In the first year of the shop's operation, the proceeds reached UAH 144 thousand.

Calculate the accounting and economic profit of the pastry shop.



1
Expert's answer
2021-02-24T15:03:50-0500
"TC=10000+24000+40000+10000+6000=90000"

accounting:


"144000-90000=54000"

economic profit


"54000-8000-(30000-6000)=22000"